# On Jiang’s wavefront sets conjecture for representations in local Arthur packets

### 13 Mar 2026 v1

@@ Abstract.
This note serves as an attempt towards the Jiang conjecture on the upper bound nilpotent orbits in the wavefront sets of representations in local Arthur packets of classical groups, which is a natural generalization of the well-known Shahidi conjecture, reflecting the relation between the structure of wavefront sets and the local Arthur parameters. Applying the character identities of local Arthur packets and the matching method of endoscopic liftings, we reduce the study of the upper bound to certain properties of the wavefront sets of the corresponding bi-torsor representations of general linear groups.
@@
The research of the first named author is partially supported by the NSF Grants DMS-1702218, DMS-1848058, and by start-up funds from the Department of Mathematics at Purdue University. The research of the second named author is partially supported by the NSF Grants DMS-1801273, DMS-2135021
@@ 1. Introduction
Then, Xu ([Xu17]) gave an algorithm to determine whether the representations in Mœglin’s construction are nonzero, and Atobe ([Ato20]) gave a refinement on the Mœglin’s construction, using the new derivatives introduced by himself and Mínguez ([AM20]), which makes it relatively easier to compute the $L$-data. Recently, Atobe ([Ato23]), Hazeltine, the first named author and Lo ([HLL22]) independently studied the intersection problem of local Arthur packets for symplectic and split odd special orthogonal groups, which is considered as a key step towards the local non-tempered Gan-Gross-Prasad problem (see [GGP20, Conjecture 7.1, Remark 7.3]).
@@
In general, characterization of the set $\mathfrak{n}^{m}(\pi)$ is still widely open, though some cases are known. Examples are representations of ${\mathrm{GL}}_{n}(F)$, irreducible subquotients for regular principal series of $G_{n}$ ([MW87]), tempered and anti-tempered unipotent representations of pure inner twists of ${\mathrm{SO}}_{2n+1}(F)$ ([Wal18, Wal20]), certain unramified representations of split connected reductive $p$-adic groups ([Oka21]), and irreducible Iwahori-spherical representations of split connected reductive $p$-adic groups with “real infinitesimal characters” ([CMO21]). It is worths to remark that recently, Tsai ([Tsa24]) constructed an example of representations of $U_{7}(\mathbb{Q}_{3})$ showing that the wavefront set $\mathfrak{p}^{m}(\pi)$ may not be a singleton.
@@
Taking the character expansion for the representation $\widetilde{\pi}_{\psi}$ of the bitorsor $\widetilde{\mathrm{GL}}_{N}(F)$ at
(1.3) $\theta_{\widehat{\mathrm{G}}_{n}}=s_{\widehat{\mathrm{G}}_{n}}\rtimes\widetilde{\theta}(N)\in\widetilde{{\mathrm{GL}}}_{N}(F),$
(see [Clo87], also see [Kon02, Theorem 3.2] and [Var17, Theorems 4.20, 4.23]), where
$s_{\widehat{\mathrm{G}}_{n}}=\begin{cases}I_{N},&\text{ when }\mathrm{G}_{n}={\mathrm{SO}}_{2n+1},\\ \begin{pmatrix}I_{n}&&\\ &1&\\ &&-I_{n}\end{pmatrix},&\text{ when }\mathrm{G}_{n}={\mathrm{Sp}}_{2n},\\ \begin{pmatrix}I_{n}&\\ &-I_{n}\end{pmatrix},&\text{ when }\mathrm{G}_{n}={\mathrm{SO}}_{2n}^{\alpha},\end{cases}$
we can define the sets $\mathfrak{n}^{m}(\widetilde{\pi}_{\psi})$ and $\mathfrak{p}^{m}(\widetilde{\pi}_{\psi})$ similarly. Note that when $\mathrm{G}_{n}={\mathrm{SO}}_{2n+1},{\mathrm{SO}}_{2n}^{\alpha}$, the connected component of the stabilizer of $\theta_{\widehat{\mathrm{G}}_{n}}$ in $\widetilde{\mathrm{GL}}_{N}(F)$ is $\widehat{\mathrm{G}}_{n}(F)$ and $\mathfrak{n}^{m}(\widetilde{\pi}_{\psi})$ consists of $F$-rational nilpotent orbits in the Lie algebra of $\widehat{\mathrm{G}}_{n}(F)$. When ${\mathrm{G}}_{n}={\mathrm{Sp}}_{2n}$, the connected component of the stabilizer of $\theta_{\widehat{\mathrm{G}}_{n}}$ in $\widetilde{\mathrm{GL}}_{2n+1}(F)$ is ${\mathrm{G}}_{n}(F)\times{\mathrm{SO}}_{1}$ and $\mathfrak{n}^{m}(\widetilde{\pi}_{\psi})$ consists of $F$-rational nilpotent orbits in the Lie algebra of ${\mathrm{G}}_{n}(F)$. Then we have the following conjecture regarding the set $\mathfrak{p}^{m}(\widetilde{\pi}_{\psi})$.
@@
[…]
where $(\underline{p}(\psi)^{t})_{\widehat{\mathrm{G}}_{n}}$ is the $\widehat{\mathrm{G}}_{n}$-collapse of the partition $\underline{p}(\psi)^{t}$ which is the largest $\widehat{\mathrm{G}}_{n}$-partition smaller than or equal to $\underline{p}(\psi)^{t}$, $(\underline{p}(\psi)^{t})^{-}$ is decreasing the smallest part of $\underline{p}(\psi)^{t}$ by $1$ and $((\underline{p}(\psi)^{t})^{-})_{{\mathrm{G}}_{n}}$ is the ${\mathrm{G}}_{n}$-collapse of $(\underline{p}(\psi)^{t})^{-}$.
@@
$\mathfrak{p}^{m}(\widetilde{\pi}_{\psi})=\begin{cases}\{(\underline{p}(\psi)^{t})_{\widehat{\mathrm{G}}_{n}}\},&\text{ when }\mathrm{G}_{n}={\mathrm{SO}}_{2n+1},{\mathrm{SO}}_{2n}^{\alpha},\\ \{((\underline{p}(\psi)^{t})^{-})_{{\mathrm{G}}_{n}}\},&\text{ when }\mathrm{G}_{n}={\mathrm{Sp}}_{2n}.\end{cases}$
@@
(1). By [CM93, Theorem 6.3.11], given a local Arthur parameter $\psi$ of $\mathrm{G}_{n}$, $(\underline{p}(\psi)^{t})_{\widehat{\mathrm{G}}_{n}}$ is always $\widehat{\mathrm{G}}_{n}$-special.
(2). Assume that $\mathrm{G}_{n}={\mathrm{SO}}_{2n+1},{\mathrm{SO}}_{2n}^{\alpha}$. When $(\underline{p}(\psi)^{t})_{\widehat{\mathrm{G}}_{n}}=\underline{p}(\psi)^{t}$, i.e., $\underline{p}(\psi)^{t}$ is already a $\widehat{\mathrm{G}}_{n}$-partition, by Theorem 1.1 and [Kon02, Theorem 4.1 (1)], $\underline{p}(\psi)^{t}\in\mathfrak{p}^{m}(\widetilde{\pi}_{\psi})$. This covers a large family of local Arthur parameters. Note that $\underline{p}(\psi)$ is automatically a $\widehat{\mathrm{G}}_{n}$-partition, then by [CM93, Proposition 6.3.7, Theorem 6.3.11], for ${\mathrm{G}}_{n}={\mathrm{SO}}_{2n+1}$, $\underline{p}(\psi)^{t}$ is a $\widehat{\mathrm{G}}_{n}$-partition if and only if $\underline{p}(\psi)$ is a $\widehat{\mathrm{G}}_{n}$-special partition; for ${\mathrm{G}}_{n}={\mathrm{SO}}^{\alpha}_{2n}$, $\underline{p}(\psi)^{t}$ is a ${\mathrm{SO}}_{2n}$-partition if and only if $\underline{p}(\psi)$ is also a ${\mathrm{Sp}}_{2n}$-partition.
When $(\underline{p}(\psi)^{t})_{\widehat{\mathrm{G}}_{n}}\neq\underline{p}(\psi)^{t}$, i.e., $\underline{p}(\psi)^{t}$ is not a $\widehat{\mathrm{G}}_{n}$-partition, Conjecture 1.3 is expected to be more complicated and is being studied by the authors.
@@
(4). Assume that ${\mathrm{G}}_{n}={\mathrm{Sp}}_{2n}$. Then
$((\underline{p}(\psi)^{t})^{-})_{{\mathrm{G}}_{n}}=((\underline{p}(\psi)^{-})_{{\mathrm{G}}_{n}})^{t}=\eta_{\mathfrak{so}_{2n+1},\mathfrak{sp}_{2n}}(\underline{p}(\psi)),$
where $\eta_{\mathfrak{so}_{2n+1},\mathfrak{sp}_{2n}}$ denotes the Barbasch-Vogan duality map from the odd orthogonal partitions to the symplectic partitions ((see [BV85] and [Ach03, Lemma 3.3])).
(5). The choices of $\theta_{\widehat{\mathrm{G}}_{n}}$ are for the purpose of taking character expansion of certain distributions at $\theta_{\widehat{\mathrm{G}}_{n}}$ and at its norm, for details, see §4.1.
@@
We prove the enhanced Shahidi conjecture assuming Conjecture 1.2, see Theorem 1.9 Part (2) below. We remark that for symplectic and split odd special orthogonal groups, Hazeltine, the first named author and Lo ([HLL22]) have proved Conjecture 1.6 without any assumption, using Atobe’s refinement on Mœglin’s construction of local Arthur packets, which is a different method from that of this paper. Hazeltine, the first named author, Lo, and Zhang ([HLLZ22]) provided a framework towards Conjecture 1.6 for general connected reductive groups, assuming the closure ordering conjecture of local Arthur packets.
@@
Let $\psi$ be a local Arthur parameter of $G_{n}$ as in (1.1), and let $\widetilde{\Pi}_{\psi}$ be the local Arthur packet attached to $\psi$. Then the followings hold.
@@
$\underline{p}_{{\mathrm{G}}_{n}}$ denotes the ${\mathrm{G}}_{n}$-collapse of $\underline{p}$, that is, the largest ${\mathrm{G}}_{n}$-partition smaller than or equal to $\underline{p}$, and $\underline{p}^{{\mathrm{G}}_{n}}$ denotes the ${\mathrm{G}}_{n}$-expansion of $\underline{p}$, that is, the smallest special ${\mathrm{G}}_{n}$-partition bigger than or equal to $\underline{p}$. Given any partition $\underline{p}=[p_{1}p_{2}\cdots p_{r}]$ with $p_{1}\geq p_{2}\geq\cdots\geq p_{r}$, let $\underline{p}^{-}=[p_{1}p_{2}\cdots(p_{r}-1)]$ and $\underline{p}^{+}=[(p_{1}+1)p_{2}\cdots p_{r}]$. For the definition of special partitions, see [CM93, Section 6.3]. We recall the definition of Barbasch-Vogan duality map from [BV85] and [Ach03] as follows.
@@
Let $\psi$ be a local Arthur parameter as in (1.1), with $\underline{p}(\psi)=[b_{1}^{a_{1}}b_{2}^{a_{2}}\cdots b_{r}^{a_{r}}]$ and $b_{1}\geq b_{2}\geq\cdots\geq b_{r}$. Assume that Conjecture 1.2 is true. Then we have the followings.
@@
    Conjecture 1.7 Part (2) is valid.
@@
    Conjecture 1.6 is true.
@@
    Let
@@
    and $n^{*}=\big\lfloor\frac{\sum_{b_{i}\text{ odd }}a_{i}}{2}\big\rfloor$. Then Conjecture 1.7 Part (3) holds for the following cases.
    1. (a)
@@
        or, $b_{i}$ are all odd,
      then (1.4) holds and thus Conjecture 1.7 Part (3) is valid.
    2. (b)
@@
      then (1.5) holds and thus Conjecture 1.7 Part (3) is valid.
    3. (c)
@@
      If all $b_{i}$ are of the same parity, then (1.6) holds and thus Conjecture 1.7 Part (3) is valid.
@@
(1) We should point out that the equations (1.4), (1.5), and (1.6) of Theorem 1.9 guarantee the existence of the member in the local Arthur packet suggested by Part (3) of Jiang’s Conjecture 1.7. In other words, this part of the conjecture is equivalent to the validity of these equations for given parameters.
(2) Conjecture 1.2 plays an important role in all parts of Theorem 1.9. In particular, for Part (3) of Theorem 1.9, we can show that there exists a member $\sigma$ in $\widetilde{\Pi}_{\psi}$ such that $\eta_{{\hat{\mathfrak{g}}_{n},\mathfrak{g}_{n}}}(\underline{p}(\psi))\in\underline{p}(\sigma)$ unconditionally, however, to show that $\eta_{{\hat{\mathfrak{g}}_{n},\mathfrak{g}_{n}}}(\underline{p}(\psi))\in\underline{p}^{m}(\sigma)$, we need Conjecture 1.2.
For Parts (1) and (2) of Theorem 1.9, we follow the idea used in [Sha90, Section 9] which is for the case of endoscopic transfer. For the Part (3) of Theorem 1.9, we follow the work of [JS04, Liu11, JL14, JL22] to construct a particular element in each local Arthur packet and study its wavefront set. We remark that when $b_{i}$ are of mixed parities, conditions (1.4) – (1.6) may not always hold, see Remark 5.7 for examples. Recently, Dr. Mathilde Gerbelli-Gauthier informed us that our results in Theorem 1.9 have possible applications on Sarnak’s density hypothesis, which quantifies how asymptotically non-tempered representations are very rare.
This paper serves as an initial step towards the deep conjecture of Jiang (Conjecture 1.7) on the wavefront sets of representations in general local Arthur packets, through the matching method of endoscopic lifting. Although the main result of this paper (Theorem 1.9 above) has a shortcoming, namely, the assumption of Conjecture 1.2, it sheds the light towards the general case of Jiang’s conjecture. By the discussions in Remark 1.4, Conjectures 1.2 and 1.3 are believed to be true and are currently being studied by the authors.
Jiang’s conjecture (Conjecture 1.7) has analogues for quasi-split unitary groups and non-quasi-split classical groups. The method used in this paper is expected to extend to these cases (as long as the local Arthur classification is carried out).
We remark that Okada ([Oka21]) has computed the wavefront set of irreducible unramified representations of split connected reductive $p$-adic groups with local Arthur parameter $\psi$ being trivial on the Weil-Deligne group. Ciubotaru, Mason-Brown, and Okada ([CMO21]) have computed the wavefront set of irreducible Iwahori-spherical representations of split connected reductive $p$-adic groups with “real infinitesimal characters”. Waldspurger ([Wal18, Wal20]) has computed the wavefront set of tempered and anti-tempered unipotent representations of pure inner twists of ${\mathrm{SO}}_{2n+1}(F)$. Combining these results on unipotent representations with the closure ordering result in [HLLZ22], jointly with Hazeltine and Lo ([HLLS24]), we proved the following theorem towards Conjecture 1.7.
@@
We remark that jointly with Hazeltine and Lo ([HLLS24]), we also proved the following very interesting reduction on Parts (1) and (2) of Conjecture 1.7.
@@
    Parts (1) and (2) of Conjecture 1.7 hold for any local Arthur parameter.
@@
    Parts (1) and (2) of Conjecture 1.7 hold for any anti-tempered local Arthur parameter.
Using similar arguments as in the proofs of Parts (1) and (2) of Theorem 1.9, assuming that Conjecture 1.7 Parts (1) and (2) are both valid, we can show that Conjecture 1.2 holds partially as follows.
Theorem 1.13.
Assume that Conjecture 1.7 Parts (1) and (2) are both valid, then Conjecture 1.2 holds partially, namely, for any
[…]
$\underline{p}\notin\mathfrak{p}^{m}(\widetilde{\pi}_{\psi})$.
@@
Let ${\mathrm{G}}_{n}={\mathrm{Sp}}_{2n},{\mathrm{SO}}_{2n+1}$. Assume that $\psi$ is trivial on $W_{F}$. Then, for any $\underline{p}\in\mathfrak{p}^{m}(\widetilde{\pi}_{\psi})$,
[…]
$\underline{p}\notin\mathfrak{p}^{m}(\widetilde{\pi}_{\psi})$.
Following is the structure of this paper. In Section 2, we recall the characterization of the local Arthur packets in [Art13, Section 2.2] and prove Theorem 1.1. In Section 3, we recall certain twisted Jacquet modules associated to nilpotent orbits, following the formulation in [GGS17]. In Section 4, we prove certain dimension identities for nilpotent orbits which are important ingredients for our main results. In Section 5, we construct a particular element in each local Arthur packet and study its wavefront set. In Section 6, we prove our main result Theorems 1.9 and 1.13. In Appendix A, joint with Alexander Hazeltine and Chi-Heng Lo, we show that representations in local Arthur packets share the same central character, which will be used in the proofs in Section 6.

### 06 May 2026 v2

# gin d [1]#1

@@ Abstract.
This paper serves as an attempt towards the Jiang conjecture on the upper bound nilpotent orbits in the wavefront sets of representations in local Arthur packets of quasi-split classical groups, which is a natural generalization of the well-known Shahidi conjecture, reflecting the relation between the structure of wavefront sets and the local Arthur parameters. Applying the character identities of local Arthur packets and a matching method, we reduce the study of the upper bound to certain properties of the wavefront sets of the corresponding bi-torsor representations of general linear groups, which is implied by a recent result of Atobe and Ciubotaru for split classical groups when the residue characteristic is large.
@@
The research of the first named author is partially supported by the NSF Grants DMS-1702218, DMS-1848058, and by the Simons Foundation: Travel Support for Mathematicians. The research of the second named author is partially supported by the NSF Grants DMS-1801273, DMS-2135021
@@ 1. Introduction
Then, Xu ([Xu17]) gave an algorithm to determine whether the representations in Mœglin’s construction are nonzero, and Atobe ([Ato20]) gave a refinement on the Mœglin’s construction, using the new derivatives introduced by himself and Mínguez ([AM20]), which makes it relatively easier to compute the $L$-data. Recently, Atobe ([Ato23]), Hazeltine, the first named author and Lo ([HLL22]) independently studied the intersection problem of local Arthur packets for symplectic and split special odd orthogonal groups, which is considered as a key step towards the local non-tempered Gan-Gross-Prasad problem (see [GGP20, Conjecture 7.1, Remark 7.3]).
@@
In general, characterization of the set $\mathfrak{n}^{m}(\pi)$ is still widely open, though some cases are known. Examples are representations of ${\mathrm{GL}}_{n}(F)$, irreducible subquotients for regular principal series of $G_{n}$ ([MW87]), tempered and anti-tempered unipotent representations of pure inner twists of ${\mathrm{SO}}_{2n+1}(F)$ ([Wal18, Wal20]), certain unramified representations of split connected reductive $p$-adic groups ([Oka21]), and irreducible Iwahori-spherical representations of split connected reductive $p$-adic groups with “real infinitesimal characters” ([CMBO21, CMBO24, CMBO25]). It is worths to remark that recently, Tsai ([Tsa24]) constructed an example of representations of $U_{7}(\mathbb{Q}_{3})$ showing that the wavefront set $\mathfrak{p}^{m}(\pi)$ may not be a singleton.
@@
Taking the character expansion for the representation $\widetilde{\pi}_{\psi}$ of the bitorsor $\widetilde{\mathrm{GL}}_{N}(F)$ at the element
(1.3) $\theta_{\widehat{\mathrm{G}}_{n}}={1}\rtimes\widetilde{\theta}(N)\in\widetilde{{\mathrm{GL}}}_{N}(F),$
(see [Clo87], also see [Kon02, Theorem 3.2] and [Var17, Theorems 4.20, 4.23]), we can define the sets $\mathfrak{n}^{m}(\widetilde{\pi}_{\psi})$ and $\mathfrak{p}^{m}(\widetilde{\pi}_{\psi})$ similarly. Let
(1.4) ${\mathrm{G}}_{n}^{\theta}:=\begin{cases}{\mathrm{SO}}_{2n+1},&\text{ when }\mathrm{G}_{n}={\mathrm{Sp}}_{2n},\\ {\mathrm{Sp}}_{2n},&\text{ when }\mathrm{G}_{n}={\mathrm{SO}}_{2n+1},{\mathrm{SO}}_{2n}^{\alpha}.\end{cases}$}
Then the connected component of the stabilizer of $\theta_{\widehat{\mathrm{G}}_{n}}$ in $\widetilde{\mathrm{GL}}_{N}(F)$ is ${\mathrm{G}}_{n}^{\theta}(F)$, and $\mathfrak{n}^{m}(\widetilde{\pi}_{\psi})$ consists of $F$-rational nilpotent orbits in the Lie algebra of ${\mathrm{G}}_{n}^{\theta}(F)$ . Then we have the following conjecture with respect to the set $\mathfrak{p}^{m}(\widetilde{\pi}_{\psi})$.
@@
$\dim_{\mathfrak{g}_{n}^{\theta}}(\underline{p})\leq\dim_{\mathfrak{g}_{n}^{\theta}}((\underline{p}(\psi)^{t})_{{\mathrm{G}}_{n}^{\theta}}),$
where $\mathfrak{g}_{n}^{\theta}$ is the Lie algebra of ${\mathrm{G}}_{n}^{\theta}$, and $(\underline{p}(\psi)^{t})_{{\mathrm{G}}_{n}^{\theta}}$ is the ${\mathrm{G}}_{n}^{\theta}$ -collapse of the partition $\underline{p}(\psi)^{t}$, namely, the largest ${\mathrm{G}}_{n}^{\theta}$-partition smaller than or equal to $\underline{p}(\psi)^{t}$.
@@
$\mathfrak{p}^{m}(\widetilde{\pi}_{\psi})={\{(}\underline{p}{(\psi)^{t})_{{\mathrm{G}}_{n}^{\theta}}\}.}$
@@
(1). When $(\underline{p}(\psi)^{t})_{{\mathrm{G}}_{n}^{\theta}}=\underline{p}(\psi)^{t}$ , i.e., $\underline{p}(\psi)^{t}$ is already a ${\mathrm{G}}_{n}^{\theta}$ -partition, by Theorem 1.1, [Kon02, Theorem 4.1 (1)], and [Var17, Lemma 5.29], for any $\underline{p}\in\mathfrak{p}^{m}(\widetilde{\pi}_{\psi})$, $\underline{p}\leq\underline{p}(\psi)^{t}_{{\mathrm{G}}_{n}^{\theta}}$, hence, Conjecture 1.2 holds. This covers a large family of local Arthur parameters. Note that $\underline{p}(\psi)$ is automatically a partition of the appropriate type for the Arthur parameter, and by [CM93, Proposition 6.3.7, Theorem 6.3.11], the condition that $\underline{p}(\psi)^{t}$ is already a ${\mathrm{G}}_{n}^{\theta}$ -partition can be checked explicitly from $\underline{p}(\psi)$ .
When $(\underline{p}(\psi)^{t})_{{\mathrm{G}}_{n}^{\theta}}\neq\underline{p}(\psi)^{t}$ , i.e., $\underline{p}(\psi)^{t}$ is not a ${\mathrm{G}}_{n}^{\theta}$ -partition, Conjecture 1.2 and 1.3 are expected to be more complicated.
@@
We prove the enhanced Shahidi conjecture assuming Conjecture 1.2 (see Theorem 1.9(3-a) below); and unconditionally when $p$ is large and $G$ is split (see Theorem 1.9(2) below). We remark that for symplectic and split special odd orthogonal groups, Hazeltine, the first named author and Lo ([HLL22]) have proved Conjecture 1.6 without any assumption, using Atobe’s refinement on Mœglin’s construction of local Arthur packets, which is a different method from that of this paper. Hazeltine, the first named author, Lo, and Zhang ([HLLZ22]) provided a framework towards Conjecture 1.6 for general connected reductive groups, assuming the closure ordering conjecture of local Arthur packets.
@@
Let $\psi$ be a local Arthur parameter of $G_{n}$ as in (1.1), and let $\widetilde{\Pi}_{\psi}$ be the local Arthur packet attached to $\psi$. Then the following hold.
@@
$\underline{p}_{{\mathrm{G}}_{n}}$ denotes the ${\mathrm{G}}_{n}$-collapse of $\underline{p}$, that is, the largest ${\mathrm{G}}_{n}$-partition smaller than or equal to $\underline{p}$, and $\underline{p}^{{\mathrm{G}}_{n}}$ denotes the ${\mathrm{G}}_{n}$-expansion of $\underline{p}$, that is, the smallest special ${\mathrm{G}}_{n}$-partition bigger than or equal to $\underline{p}$. Given any partition $\underline{p}=[p_{1}p_{2}\cdots p_{r}]$ with $p_{1}\geq p_{2}\geq\cdots\geq p_{r}$, let $\underline{p}^{-}=[p_{1}p_{2}\cdots(p_{r}-1)]$ and $\underline{p}^{+}=[(p_{1}+1)p_{2}\cdots p_{r}]$. For the definition of special partitions, see [CM93, Section 6.3]. We recall the definition of the Barbasch-Vogan duality map from [BV85] and [Ach03] as follows.
@@
Let $\psi$ be a local Arthur parameter as in (1.1), with $\underline{p}(\psi)=[b_{1}^{a_{1}}b_{2}^{a_{2}}\cdots b_{r}^{a_{r}}]$ and $b_{1}\geq b_{2}\geq\cdots\geq b_{r}$. Then, we have the following.
@@
    (Theorem 6.4). Conjecture 1.2 is true if and only if for any $\pi\in\widetilde{\Pi}_{\psi}$, $\underline{p}\in\mathfrak{p}^{m}(\pi)$,
    $\dim_{{\mathfrak{g}}_{n}}(\underline{p})\leq\dim_{{\mathfrak{g}}_{n}}(\eta_{{\hat{\mathfrak{g}}_{n},\mathfrak{g}_{n}}}(\underline{p}(\psi))).$
    In particular, the latter implies that Conjecture 1.7(2) is valid.
@@
    For split ${\mathrm{G}}_{n}$, Conjecture 1.2 holds when $p$ is large as follows:
    1. (a)
      $p>6n+3$, when ${\mathrm{G}}_{n}={\mathrm{SO}}_{2n+1}$;
    2. (b)
      $p>6n$, when ${\mathrm{G}}_{n}={\mathrm{Sp}}_{2n}$;
    3. (c)
      $p>6n$, when ${\mathrm{G}}_{n}={\mathrm{SO}}_{2n}$.
@@
    Assume that Conjecture 1.2 is true. Then we have the following.
    1. (a)
      (Theorem 6.5). Conjecture 1.6 is true.
    2. (b)
      (Theorem 6.6). Let
@@
      and $n^{*}=\big\lfloor\frac{\sum_{b_{i}\text{ odd }}a_{i}}{2}\big\rfloor$. Then Conjecture 1.7(3) holds for the following cases.
      1. (i)
@@
        then (1.6) holds and thus Conjecture 1.7(3) is valid.
      3. (iii)
@@
        If all $b_{i}$ are of the same parity, then (1.7) holds and thus Conjecture 1.7(3) is valid.
@@
  1. (1)
    Theorem 1.9(1) is proved by applying a generalization of the matching method used in [Sha90, Section 9] to the endoscopic character identity in [Art13].
  2. (2)
    For split ${\mathrm{G}}_{n}$, applying the equivalence in Theorem 1.9(1), Theorem 1.9(2) is directly implied by [AC26, Theorem 1.4(2)]. Indeed, [AC26] claims a proof of Conjecture 1.7 for split $G_{n}$ under certain hypothesis, applying similar matching methods.
  3. (3)
    For Theorem 1.9(3-b), we explicitly construct a member $\sigma$ in $\widetilde{\Pi}_{\psi}$ such that $\eta_{{\hat{\mathfrak{g}}_{n},\mathfrak{g}_{n}}}(\underline{p}(\psi))\in\underline{p}(\sigma)$, however, to show that $\eta_{{\hat{\mathfrak{g}}_{n},\mathfrak{g}_{n}}}(\underline{p}(\psi))\in\underline{p}^{m}(\sigma)$, we need Conjecture 1.2. The equalities (1.5) - (1.7) guarantee that $\eta_{{\hat{\mathfrak{g}}_{n},\mathfrak{g}_{n}}}(\underline{p}(\psi))\in\underline{p}(\sigma)$. When $b_{i}$ are of mixed parities, then the equalities (1.5) – (1.7) may not always hold, see Remark 5.7 for examples.
This paper serves as an initial step towards the deep conjecture of Jiang (Conjecture 1.7) on the wavefront sets of representations in general local Arthur packets, through the endoscopic character identities. Jiang’s conjecture (Conjecture 1.7) has analogs for quasi-split unitary groups and non-quasi-split classical groups. The method used in this paper is expected to extend to these cases (as long as the local Arthur classification is carried out).
We remark that Okada ([Oka21]) has computed the wavefront set of irreducible unramified representations of split connected reductive $p$-adic groups with local Arthur parameter $\psi$ being trivial on the Weil-Deligne group. Ciubotaru, Mason-Brown, and Okada ([CMBO24, CMBO25]) have computed the wavefront set of irreducible Iwahori-spherical representations of split connected reductive $p$-adic groups with “real infinitesimal characters”. Waldspurger ([Wal18, Wal20]) has computed the wavefront set of tempered and anti-tempered unipotent representations of pure inner twists of ${\mathrm{SO}}_{2n+1}(F)$. Combining these results on unipotent representations with the closure ordering result in [HLLZ22], jointly with Hazeltine and Lo ([HLLS24]), we proved the following theorem towards Conjecture 1.7.
@@
We remark that, together with Hazeltine and Lo ([HLLS24]), we also proved the following very interesting reduction in Conjecture 1.7(1-2).
@@
    Conjecture 1.7(1-2) holds for any local Arthur parameter.
@@
    Conjecture 1.7(1-2) holds for any anti-tempered local Arthur parameter.
  3. (3)
    Conjecture 1.7(1-2) holds for any anti-discrete local Arthur parameter.
@@
Let ${\mathrm{G}}_{n}={\mathrm{Sp}}_{2n},{\mathrm{SO}}_{2n+1}$. Assume that $\psi$ is trivial on $W_{F}$. Then Conjecture 1.2 holds.
The following is the structure of this paper. In Section 2, we recall the characterization of the local Arthur packets in [Art13, Section 2.2] and prove Theorem 1.1. In Section 3, we recall certain twisted Jacquet modules associated to nilpotent orbits, following the formulation in [GGS17]. In Section 4, we prove certain dimension identities for nilpotent orbits, which are important ingredients for our main results. In Section 5, we construct a particular element in each local Arthur packet and study its wavefront set. In Section 6, we prove our main result Theorems 1.9. In Appendix A, together with Alexander Hazeltine and Chi-Heng Lo, we show that all representations in any local Arthur packet share the same central character, which has its own interests.